Here is another free question. Radius of a shpere; the raduis of a sphere is given by the formula r = (0.75V/pi sign)1/3, where V is its volume. Find the radius of a sphere tank that has a volume of 32pi/3 cubic meters
Answer by mananth(16946)

r = (0.75V/pi)^1/3
V= 32*pi/3
r= (0.75*32*pi/3*pi)^1/3
r=(0.75/3)^1/3
0.083m
...
m.ananth@hotmail.ca